36.53 mL of a 0.223 M solution of weak acid HA are titrated with 0.2 M NaOH. What is the pH of the solution after 8.25 mL of the NaOH have been added?

The Ka for the weak acid is 0.0000077101.
Answer #1

Given:

M(HA) = 0.223 M

V(HA) = 36.53 mL

M(NaOH) = 0.2 M

V(NaOH) = 8.25 mL

mol(HA) = M(HA) * V(HA)

mol(HA) = 0.223 M * 36.53 mL = 8.1462 mmol

mol(NaOH) = M(NaOH) * V(NaOH)

mol(NaOH) = 0.2 M * 8.25 mL = 1.65 mmol

We have:

mol(HA) = 8.1462 mmol

mol(NaOH) = 1.65 mmol

1.65 mmol of both will react

excess HA remaining = 6.4962 mmol

Volume of Solution = 36.53 + 8.25 = 44.78 mL

[HA] = 6.4962 mmol/44.78 mL = 0.1451M

[A-] = 1.65/44.78 = 0.0368M

They form acidic buffer

acid is HA

conjugate base is A-

Ka = 7.71*10^-6

pKa = - log (Ka)

= - log(7.71*10^-6)

= 5.113

use:

pH = pKa + log {[conjugate base]/[acid]}

= 5.113+ log {3.685*10^-2/0.1451}

= 4.518

Answer: 4.52